# Fractal and Regular Geometry of Deep Neural Networks

### v1

@@ 4.1   Hausdorff dimension and measure
We briefly recall the definition of Hausdorff dimension and some of its well known properties (see [27] for more details).
@@
$\dim_{\rm H}(F)=\sup_{s\geq 0}\left\{\mathcal{H}^{s}(F)=\infty\right\}\ ,$
@@
$\dim_{\rm H}(F)=\inf_{s\geq 0}\left\{\mathcal{H}^{s}(F)=0\right\}\ .$
The following theorem, whose proof can be found in [27], Theorem 4.13, allows us to obtain a lower bound for the Hausdorff dimension of a set. The proof is given only for subsets of $\mathbb{R}^{n}$, but it can be easily generalized to any metric space.
@@
The following proposition generalizes to Borel subsets of $\mathbb{S}^{d}\times\mathbb{R}$ the result proved for Borel subsets of $\mathbb{R}^{2}$ in [27], Proposition 7.9. The proof follows the one presented in [27], with slight modifications. Once this result is established, we obtain Corollary 4.3 in the same way as in [27], Corollary 7.10.
@@
Let $F$ be a Borel subset of $\mathbb{S}^{d}\times\mathbb{R}$. Then, for every $11\leq s\leq d+$ 1 we have
@@
Let $\delta>0$, and let $(U_{i})_{i\in I}$ be a $\delta$-cover of $F$ such that
@@
For each $i$, there exists a $(d+1)$-dimensional square of side-length $\mathrm{diam}(U_{i})$ and an interval $R_{i}\subseteq\mathbb{R}$ of length $\mathrm{diam}(U_{i})$ such that
$U_{i}\subseteq S_{i}\times R_{i}\ .$
It is easy to see that
$U_{i}\cap L_{x}\subseteq\begin{cases}\varnothing&\text{ if }(t,x)\not\in U_{i}\text{ for any $t\in U_{i}$ }\\ S_{i}\times\left\{x\right\}&\text{ otherwise}\end{cases}\ ,$
and that $(U_{i}\cap L_{x})_{i\in I}$ is a $\delta$-cover of $F\cap L_{x}$. So
$\mathcal{H}^{s-1}_{\delta}(F\cap L_{x})\leq$ $\sum_{i\in I}\mathrm{diam}(S_{i}\cap L_{x})^{s-1}=\sum_{i\in I}\mathrm{diam}(U_{i})^{s-1}$
$=$ $\sum_{i\in I}\mathrm{diam}(U_{i})^{s-d-2}\mathrm{diam}(U_{i})^{d+1}$
$=$ $\sum_{i\in I}\mathrm{diam}(U_{i})^{s-d-2}\int_{R_{i}}\mathbb{I}_{U_{i}\times R_{i}}((t,x))\operatorname{dVol}(t)\ .$
Hence,
$\int_{\mathbb{R}}\mathcal{H}^{s-1}_{\delta}(F\cap L_{x})$ $\leq\sum_{i\in I}\mathrm{diam}(U_{i})^{s-d-2}\int_{\mathbb{R}}\int_{R_{i}}\mathbb{I}_{S_{i}\times R_{i}}((t,x))\operatorname{dVol}(t)\mathrm{d}x$
$\leq\sum_{i\in I}\mathrm{diam}(U_{i})^{s-d-2}\mathrm{diam}(U_{i})^{d+2}\leq\mathcal{H}^{s}_{\delta}(F)+\varepsilon\ .$
@@
$\int_{\mathbb{R}}\mathcal{H}^{s-1}(F\cap L_{x})\leq\mathcal{H}^{s}_{\delta}(F)$
and the claim follows by letting $\delta\to 0$. ∎
@@ 4.2   Technical tools
In this section we give two technical tools which are useful to compute the Hausdorff dimension. The first result generalizes [37], Theorem 2 from dimension $2$ to any dimension. Its proof is rather technical and follows the argument in [37] (see Appendix B for more details). The second tool that we will use in the proof of Theorem 3.9 is the construction of a $\delta$-net on the sphere; in our second technical lemma we prove a bound for its cardinality.
@@
Let $T$ be an isotropic Gaussian random field with CRI $\beta\in(0,1)$. Then there exists a constant $K_{1}>0$ such that, with probability one,
$\lim_{\varepsilon\to 0}\sup_{x,y\in\mathbb{S}^{d}\atop{d_{\mathbb{S}^{d}}(x,y)\leq\varepsilon}}\frac{|T(x)-T(y)|}{d_{\mathbb{S}^{d}}(x,y)^{\beta}\sqrt{\ln d_{\mathbb{S}^{d}}(x,y)}}=K_{1}\ ,$
where $d_{\mathbb{S}^{d}}$ denotes the geodesic distance on the sphere, i.e. $d_{\mathbb{S}^{d}}(x,y)=\arccos(\langle x,y\rangle)$.
Let $S$ be a subset of $\mathbb{S}^{d}$, and let $\delta>0$. We say that $S$ is $\delta$-separated if $d_{\mathbb{S}^{d}}(x,y)>\delta$ for all $x,y\in S$ with $x\neq y$; furthermore we say that $S$ is maximal if $S\cup\left\{x\right\}$ is not $\delta$-separated anymore, for any $x\in\mathbb{S}^{d}$.
@@
For all $\delta>0$, there exists a maximal $\delta$-separated set $S_{\delta}$. Moreover, for $\delta\to 0$ we have
@@
where $C_{d}=\frac{d}{2}\operatorname{B}(d/2,1/2)$ and $\operatorname{B}$ denotes the Beta function.
@@
We split the proof in three parts.
  • Let $\mathcal{S}$ be the set of all $\delta$-separated sets. Let $x\in\mathbb{S}^{d}$; then $\left\{x\right\}$ is $\delta$-separated. Hence $(\mathcal{S},\subseteq)$ is a non-empty poset. If we prove that every chain in $\mathcal{S}$ has an upper bound in $\mathcal{S}$, the first part of the claim follows by Zorn’s lemma. Let $\mathcal{C}\subseteq\mathcal{S}$ be a chain and $U=\bigcup\mathcal{C}$; then $U\in\mathcal{S}$. Indeed, for all $x,y\in U$ there exist $C,D\in\mathcal{C}$ such that $x\in C$ and $y\in D$. Since $\mathcal{C}$ is a chain, we can assume without less of generality that $C\subseteq D$. Hence $x,y\in D$. Since $D\in\mathcal{S}$, $D$ is $\delta$-separated, and we have $d_{\mathbb{S}^{d}}(x,y)>\delta$.
  • We now prove the lower bound for the cardinality of $S$. Let $B_{\delta}(x)=\left\{y\in\mathbb{S}^{d}\;|\;d_{\mathbb{S}^{d}}(x,y)\leq\delta\right\}$. Then, by the maximality of $S$,
@@
    where $x$ is an arbitrary point on $\mathbb{S}^{d}$ and $|S|$ denotes the cardinality of $S$. Now, using the spherical coordinates we have
@@
    For the numerator, using the change of variable $\cos^{2}(\theta)=t$ we obtain
@@ 4.3   The CRI of deep neural networks
In this section we prove that the fractal class is stable under the depth $L$ and we show how the Hausdorff dimension increases with $L$.
@@
For $L=1$, the claim coincides with (3.1). For $L\geq 1$ we have
@@
To compute the Hausdorff dimension of $\Gamma_{T}$, we prove an upper bound by exhibiting an explicit cover, and a matching lower bound using a standard argument based on the potential method (see Theorem 4.1).
In view of Proposition 4.4, for any $\delta>0$ small enough and any $x,y\in\mathbb{S}^{d}$ with $d_{\mathbb{S}^{d}}(x,y)\leq\delta$,
@@
where $B_{\delta}((x,s))$ denotes the ball of center $(x,s)\in\mathbb{S}^{d}\times\mathbb{R}$ and radius $\delta$ with respect to the metric
@@
Let $s_{\varepsilon}=d+1-\beta+\varepsilon$ for some $\varepsilon>0$. Then, using Lemma 4.5 we have
@@
where $C^{\prime}$ is a constant depending on $d,\varepsilon,\beta$. Since $10<\beta<$ 1, we have $\mathcal{H}^{s_{\varepsilon}}(\Gamma_{T})=0$. Since $\varepsilon$ is arbitrary, we get the upper bound
@@
To prove the lower bound, we use the potential method (cf. Theorem 4.1). Let $\mu$ be the mass distribution on $\Gamma_{T}$ defined as
@@
$A_{s}=\mathbb{E}\left[\iint_{\Gamma_{T}^{2}}d(z,w)^{-s}\mathrm{d}\mu(z)\mathrm{d}\mu(w)\right]=\iint_{(\mathbb{S}^{d})^{2}}\mathbb{E}\left[\left(|T(x)-T(y)|\vee d_{\mathbb{S}^{d}}(x,y)\right)^{-s}\mathrm{d}x\mathrm{d}y\right]\ .$
If $x,y\in\mathbb{S}^{d}$, we denote with $\theta_{x,y}$ the angle beetween $x$ and $y$. By Proposition 4.4, there exists $\theta_{0}>0$ such that, for any $x,y$ such that $\theta_{x,y}<\theta_{0}$, we have
$|T(x)-T(y)|\vee\theta_{x,y}\geq|T(x)-T(y)|\geq\frac{K_{1}}{2}\theta_{x,y}^{\beta}\sqrt{|\ln(\theta_{x,y})|}$ (4.3)
@@
Combining (4.3) and (4.4), using spherical coordinates, since $-s\beta+d-1>-1$ we get
$A_{s}\leq\frac{\omega_{d}K_{1}}{2}\int_{0}^{\theta_{0}}\theta^{-s\beta}|\ln(\theta)|^{-s/2}\theta^{d-1}\mathrm{d}\theta+\int_{\theta_{0}}^{\pi}\theta^{-s+d-1}\mathrm{d}\theta<\infty\ .$
@@
almost surely in $\Omega$. By Theorem 4.1 we finally obtain
@@
To obtain the lower bound, one can proceed as in the previous section, replacing $\mu$ with $\nu$, an absolutely continuous measure with respect to $\mu$ such that, for any $\nu$ measurable function $f$,
$\int_{\Gamma_{T}\cap L_{u}}f(x,t)\mathrm{d}\nu((x,t))=\int_{\Gamma_{T}\cap L_{u}}f(x,t)\mathbbm{1}_{\left\{u\right\}}(t)\mathrm{d}\mu((x,t))\ .$
Let $s=d-\beta$. Then we have
$\iint_{(\Gamma_{T}\cap L_{u})^{2}}d(z,w)^{-s}\mathrm{d}\nu(z)\mathrm{d}\nu(w)$
[…]
$\leq$ $\iint_{\mathbb{S}^{d}\times\mathbb{S}^{d}}d_{\mathbb{S}^{d}}(x,y)^{-s}\mathrm{d}x\mathrm{d}y=P_{d}\int_{0}^{\pi}\theta^{-s+d-1}\mathrm{d}\theta\ ,$
for some constant $P_{d}$. Since $-s+d-1=-1+\beta>-1$ we get
$\iint_{(\Gamma_{T}\cap L_{u})^{2}}d(z,w)^{-s}\mathrm{d}\nu(z)\mathrm{d}\nu(w)<\infty\ .$
Using Theorem 4.1 we obtain
$\dim_{\rm H}(\Gamma_{T}\cap L_{u})\geq d-\beta$
for every $u\in\mathbb{R}$. For the upper bound one can use Corollary 4.3, which complets the proof of Theorem 3.9.
5   Proof of Proposition 3.14: Kac-Rice class implies $C^{1}$
@@
$\varphi_{N}\left(\begin{pmatrix}x_{1},\dots,x_{d+1}\end{pmatrix}\right)=\begin{pmatrix}\frac{x_{1}}{1-x_{d+1}},\frac{x_{2}}{1-x_{d+1}},\dots,\frac{x_{d}}{1-x_{d+1}}\end{pmatrix}\ ,$
$\varphi_{S}\left(\begin{pmatrix}x_{1},\dots,x_{d+1}\end{pmatrix}\right)=\begin{pmatrix}-\frac{x_{1}}{1+x_{d+1}},\frac{x_{2}}{1+x_{d+1}},\dots,\frac{x_{d}}{1+x_{d+1}}\end{pmatrix}\ .$
@@
$X_{n}\xrightarrow{L^{2}}X\quad\Longleftrightarrow\quad\lim_{n\to+\infty}\lim_{m\to+\infty}\mathbb{E}[X_{n}X_{m}]<\infty\ .$

### v2

@@ 4.1 Hausdorff dimension and measure
We briefly recall the definition of Hausdorff dimension and some of its well-known properties (see [30] for more details).
@@
$\dim_{\rm H}(Y)=\sup_{s\geq 0}\left\{\mathcal{H}^{s}(Y)=\infty\right\}\ ,$
@@
$\dim_{\rm H}(Y)=\inf_{s\geq 0}\left\{\mathcal{H}^{s}(Y)=0\right\}\ .$
The following theorem, whose proof can be found in [30], Theorem 4.13, allows us to obtain a lower bound for the Hausdorff dimension of a set.
@@
Remark 4.2.
The proof of the previous theorem is given only for subsets of $\mathbb{R}^{n}$. In the following we use this theorem also for $F\subseteq\mathbb{S}^{d}\times\mathbb{R}$, using the fact that the geodesic distance on $\mathbb{S}^{d}$ and the restriction on $\mathbb{S}^{d}$ of the Euclidian metric are equivalent. Furthermore, since for two metric spaces $(X,d_{X})$ and $(Y,d_{Y})$ the metrics
$\begin{aligned} &d_{2}((x,y),(x^{\prime},y^{\prime}))=\sqrt{d_{X}(x,x^{\prime})^{2}+d_{Y}(y,y^{\prime})^{2}}\,,\\ &d_{\infty}((x,y),(x^{\prime},y^{\prime}))=d_{X}(x,x^{\prime})\vee d_{Y}(y,y^{\prime})\end{aligned}\qquad x,x^{\prime}\in X,\;y,y^{\prime}\in Y$
are equivalent, we will from now on use $d_{\infty}$ on product spaces unless specified otherwise.
The following proposition generalizes to Borel subsets of $\mathbb{S}^{d}\times\mathbb{R}$ the result proved for Borel subsets of $\mathbb{R}^{2}$ in [30], Proposition 7.9. The proof follows the one presented in [30], with slight modifications. Once this result is established, we obtain Corollary˜4.4 in the same way as in [30], Corollary 7.10.
@@
Let $F$ be a Borel subset of $\mathbb{S}^{d}\times\mathbb{R}$. Then, for every $1\leq s\leq d+2$ we have
@@
Following Remark˜4.2, we consider $\mathbb{S}^{d}\times\mathbb{R}$ seen as a subset of $\mathbb{R}^{d+2}$ with the infinity metric $d_{\infty}$. Let $\delta>0$. For any $\varepsilon>0$, let $(U_{i})_{i\in I}$ be a $\delta$-cover of $F$ such that
@@
For every $j=1,\dots,d+2$, let $\pi_{j}:\,\mathbb{R}^{d+2}\to\mathbb{R}$ be the projection on the $j$-th component. Since we are using $d_{\infty}$, for any $i\in I$ and $j=1,\dots,d+2$,
$\mathrm{diam}(\pi_{j}(U_{i}))\leq\mathrm{diam}(U_{i})\,.$
It follows that, for every $i\in I$ and $j=1,\dots,d+2$, there exists $Q_{j}^{i}\subseteq\mathbb{R}$ such that $\pi_{j}(U_{i})\subseteq Q_{j}^{i}$ and $\mathrm{diam}(Q_{j}^{i})=\mathrm{diam}(U_{i})$. Then
$U_{i}\subseteq\prod_{j=1}^{d+2}Q_{j}^{i}=S_{i}\ .$
We note that $\mathrm{diam}(S_{i})=\mathrm{diam}(U_{i})$ and hence $(S_{i}\cap L_{x})_{i\in I}$ is a $\delta$-cover of $F\cap L_{x}$. Let
$I_{x}=\{i\in I\;|\;x\in Q_{d+2}^{i}\}\,.$
Then
$S_{i}\cap L_{x}=\begin{cases}\prod_{j=1}^{d+1}Q_{j}^{i}\times\left\{x\right\}&\text{ if }i\in I_{x}\\ \emptyset&\text{ otherwise}\end{cases}$
and in particular
$\mathrm{diam}(S_{i}\cap L_{x})=\begin{cases}\mathrm{diam}(U_{i})&\text{ if }i\in I_{x}\\ 0&\text{ otherwise}\end{cases}\,.$
When $I_{x}\neq\emptyset$, one has
$\mathcal{H}^{s-1}_{\delta}(F\cap L_{x})\leq$ $\sum_{i\in I}\mathrm{diam}(S_{i}\cap L_{x})^{s-1}=\sum_{i\in I_{x}}\mathrm{diam}(S_{i}\cap L_{x})^{s-1}$
$=$ $\sum_{i\in I_{x}}\mathrm{diam}(S_{i}\cap L_{x})^{s-d-2}\mathrm{diam}(S_{i}\cap L_{x})^{d+1}$
$=$ $\sum_{i\in I_{x}}\mathrm{diam}(U_{i})^{s-d-2}\int_{\mathbb{R}^{d+1}}\mathbb{I}_{S_{i}\cap L_{x}}((t,x))\mathrm{d}t$
$=$ $\sum_{i\in I}\mathrm{diam}(U_{i})^{s-d-2}\int_{\mathbb{R}^{d+1}}\mathbb{I}_{S_{i}\cap L_{x}}((t,x))\mathrm{d}t\ .$
We note that the previous inequality holds also when $I_{x}=\emptyset$, since both left-and right-hand sides are zero. Thus,
$\int_{\mathbb{R}}\mathcal{H}^{s-1}_{\delta}(F\cap L_{x})\mathrm{d}x$ $\leq\sum_{i\in I}\mathrm{diam}(U_{i})^{s-d-2}\int_{\mathbb{R}}\int_{\mathbb{R}^{d+1}}\mathbb{I}_{S_{i}\cap L_{x}}((t,x))\mathrm{d}t\mathrm{d}x$
$=\sum_{i\in I}\mathrm{diam}(U_{i})^{s-d-2}\mathrm{diam}(U_{i})^{d+2}\leq\mathcal{H}^{s}_{\delta}(F)+\varepsilon\ .$
@@
$\int_{\mathbb{R}}\mathcal{H}^{s-1}(F\cap L_{x})\mathrm{d}\leq\mathcal{H}^{s}_{\delta}(F)$
and the claim follows by letting $\delta\to 0$.
@@ 4.2 Technical tools
In this section we give two technical tools which are useful to compute the Hausdorff dimension. The first result generalizes [40], Theorem 2 from dimension $2$ to any dimension. Its proof is rather technical and follows the argument in [40] (see Appendix˜B for more details). The second tool that we will use in the proof of Theorem˜3.11 is the construction of a $\delta$-net on the sphere; in our second technical lemma, we prove a bound for its cardinality.
@@
Let $T$ be an isotropic Gaussian random field with CRI $\beta\in(0,1]$. Then there exists a constant $K>0$ such that, with probability one,
$\lim_{\varepsilon\to 0}\sup_{x,y\in\mathbb{S}^{d}\atop{d_{\mathbb{S}^{d}}(x,y)\leq\varepsilon}}\frac{|T(x)-T(y)|}{d_{\mathbb{S}^{d}}(x,y)^{\beta}\sqrt{|\ln d_{\mathbb{S}^{d}}(x,y)|}}\leq K\ ,$
where $d_{\mathbb{S}^{d}}$ denotes the geodesic distance on the sphere, i.e. $d_{\mathbb{S}^{d}}(x,y)=\arccos(\langle x,y\rangle)$. If $\beta<1$, then the display above holds as an equality.
Let $S$ be a subset of $\mathbb{S}^{d}$, and let $\delta>0$. We say that $S$ is $\delta$-separated if $d_{\mathbb{S}^{d}}(x,y)>\delta$ for all $x,y\in S$ with $x\neq y$; furthermore, we say that $S$ is maximal if $S\cup\left\{x\right\}$ is not $\delta$-separated anymore, for any $x\in\mathbb{S}^{d}$.
@@
For all $\delta>0$, there exists a maximal $\delta$-separated set $S_{\delta}$ of $\mathbb{S}^{d}$. Moreover, for $\delta\to 0$ we have
@@
where $C_{d}=\frac{d}{2}\operatorname{B}(d/2,1/2)$ with $\operatorname{B}$ denoting the Beta function, and $|S_{\delta}|$ is the cardinality of $S_{\delta}$.
@@
We split the proof into three parts.
  • Let $\mathcal{S}$ be the set of all $\delta$-separated sets. Let $x\in\mathbb{S}^{d}$; then $\left\{x\right\}$ is $\delta$-separated. Hence $(\mathcal{S},\subseteq)$ is a non-empty poset. If we prove that every chain in $\mathcal{S}$ has an upper bound in $\mathcal{S}$, the first part of the claim follows by Zorn’s lemma. Let $\mathcal{C}\subseteq\mathcal{S}$ be a chain and $U=\bigcup\mathcal{C}$; then $U\in\mathcal{S}$. Indeed, for all $x,y\in U$ there exist $C,D\in\mathcal{C}$ such that $x\in C$ and $y\in D$. Since $\mathcal{C}$ is a chain, we can assume without loss of generality that $C\subseteq D$. Hence $x,y\in D$. Since $\in\mathcal{S}, DD$ is $\delta$-separated, and we have $d_{\mathbb{S}^{d}}(x,y)>\delta$.
  • Let $S=S_{\delta}$ be a maximal $\delta$-separated set. We now prove the lower bound for the cardinality of $S$. Let $B_{\delta}(x)=\left\{y\in\mathbb{S}^{d}\;|\;d_{\mathbb{S}^{d}}(x,y)\leq\delta\right\}$. Then, by the maximality of $S$,
@@
    where $x$ is an arbitrary point on $\mathbb{S}^{d}$. Now, using the spherical coordinates we have
@@
    For the numerator, using the change of variable $\cos(\theta)^{2}=t$ we obtain
@@ 4.3 The CRI of deep neural networks
In this section, we prove that the fractal class is stable under the depth $L$ and we show how the Hausdorff dimension increases with $L$.
@@
For $L=1$, the claim coincides with (3.1), since in the case $\beta\in(0,1)$ one has $p_{1}\equiv 1$, bacause $\kappa(1)=1$. For $L\geq 1$ we have
@@
To compute the Hausdorff dimension of $\Gamma_{T}$, we prove an upper bound by exhibiting an explicit cover, and a matching lower bound using a standard argument based on the potential method (see Theorem˜4.1).
In view of Proposition˜4.5, for any $\delta>0$ small enough and any $x,y\in\mathbb{S}^{d}$ with $d_{\mathbb{S}^{d}}(x,y)\leq\delta$,
@@
where $B_{\delta}((x,s))$ denotes the ball of centre $(x,s)\in\mathbb{S}^{d}\times\mathbb{R}$ and radius $\delta$ with respect to the metric
@@
Let $s_{\varepsilon}=d+1-\beta+\varepsilon$ for some $\varepsilon>0$. Then, using Lemma˜4.6 and the definition of $N$ we have
@@
where $C^{\prime}$ is a constant depending on $d,\varepsilon,\beta$. Since $0<\beta<1$, we have $\mathcal{H}^{s_{\varepsilon}}(\Gamma_{T})=0$. Since $\varepsilon$ is arbitrary, we get the upper bound
@@
To prove the lower bound, we use the potential method (cfr. Theorem˜4.1). Let $\mu$ be the mass distribution on $\Gamma_{T}$ defined as
@@
$A_{s}=\mathbb{E}\left[\iint_{\Gamma_{T}^{2}}d(z,w)^{-s}\mathrm{d}\mu(z)\mathrm{d}\mu(w)\right]=\iint_{(\mathbb{S}^{d})^{2}}\mathbb{E}\left[\left(|T(x)-T(y)|\vee d_{\mathbb{S}^{d}}(x,y)\right)^{-s}\right]\mathrm{d}x\mathrm{d}y\ .$
If $x,y\in\mathbb{S}^{d}$, we denote with $\theta_{x,y}$ the angle between $x$ and $y$. We split the proof in two parts, one for $\beta\in(0,1)$ and one for $\beta=1$.
$\beta\in(0,1)$
By Proposition˜4.5, there exists $\theta_{0}>0$ such that, for any $x,y$ such that $\theta_{x,y}<\theta_{0}$, we have
$|T(x)-T(y)|\vee\theta_{x,y}\geq|T(x)-T(y)|\geq\frac{K}{2}\theta_{x,y}^{\beta}\sqrt{|\ln(\theta_{x,y})|}$ (4.3)
@@
Combining (4.3) and (4.4), using spherical coordinates, since $s\beta+d-1>-1$ we get
$A_{s}\leq\frac{\omega_{d}K}{2}\int_{0}^{\theta_{0}}\theta^{-s\beta}|\ln(\theta)|^{-s/2}\theta^{d-1}\mathrm{d}\theta+\omega_{d}\int_{\theta_{0}}^{\pi}\theta^{-s+d-1}\mathrm{d}\theta<\infty\ .$
@@
almost surely in $\Omega$. By Theorem˜4.1 we finally obtain
@@
$\beta=1$
In view of (4.4), for every $\varepsilon\in(0,d)$ we have
$A_{d-\varepsilon}\leq\int_{0}^{\pi}\frac{\mathrm{d}\theta}{\theta^{1-\varepsilon}}<\infty\ .$
Using again Theorem˜4.1 we have
$\dim_{\rm H}(\Gamma_{T})\geq d-\varepsilon\ .$
Thus
$d-\varepsilon\leq\dim_{\rm H}(\Gamma_{T})\leq d$
for every $\varepsilon>0$, whence the claim.
@@
To obtain the lower bound we rely again on the potential method. The argument is somewhat technical and closely follows the approach in [2]. For completeness, we provide the full derivation in Appendix˜C. More precisely, we show that, for any fixed $u\in\mathbb{R}$, there exists a (random) measure $\mu$ such that, for every $\gamma<d-\beta$,
$\int_{T^{-1}(u)\times T^{-1}(u)}\frac{1}{\|z-w\|^{\gamma}}\,\mathrm{d}\mu(z)\,\mathrm{d}\mu(w)<\infty\qquad\text{a.s.}$
and that, with positive probability, $\mu$ is a mass distribution supported on $T^{-1}(u)$. Applying Theorem˜4.1, we conclude that
$\dim_{\mathrm{H}}(\Gamma_{T}\cap L_{u})\;\geq\;d-\beta\,.$
For the upper bound one can use Corollary˜4.4, which completes the proof of Theorem˜3.11.
5  Proof of Proposition˜3.16: Kac-Rice class implies $C^{1}$
@@
$\varphi_{N}\left(\begin{pmatrix}x_{1},\dots,x_{d+1}\end{pmatrix}\right)=\begin{pmatrix}\frac{x_{1}}{1-x_{d+1}},\frac{x_{2}}{1-x_{d+1}},\dots,\frac{x_{d}}{1-x_{d+1}}\end{pmatrix}\$,
$\varphi_{S}\left(\begin{pmatrix}x_{1},\dots,x_{d+1}\end{pmatrix}\right)=\begin{pmatrix}-\frac{x_{1}}{1+x_{d+1}},\frac{x_{2}}{1+x_{d+1}},\dots,\frac{x_{d}}{1+x_{d+1}}\end{pmatrix}\$.
@@
$X_{n}\xrightarrow{L^{2}}X\quad\Longleftrightarrow\quad\lim_{n,m\to+\infty}\mathbb{E}[X_{n}X_{m}]<\infty\ .$